OpenStudy (gracie007):

So I need to solve log(x+9)=2, the answer is 81 but I have no idea how to get to that. Help please?

OpenStudy (hafeda):

u sure the answer is 81 or 91. cuz i am getting 91

OpenStudy (dustin_whitelock):

ANS should be 91.

OpenStudy (gracie007):

it's 91, i had a typo

OpenStudy (hafeda):

so do u know how to solve it, or do i post a solution

OpenStudy (gracie007):

i have no idea how to solve it.

OpenStudy (hafeda):

log(x+9) = 2 the base of the log is 10, because a log with no base indicated it is always 10. Method 1 : you can bring the base of the log which is 10 to the other side of the equation which will look like as follow: x+9 = 10^2 (log is gone ) x+9 = 100 ( Subtract 9 from both side ) x= 100 - 9 x= 91. Method 2: You can rise both side of the equation by 10, which will look as follow: 10^log(x+9) = 10^2 ( Now we can use the inverse of log aka antilog which is 10^log n ) x+9 = 100 (by this logic the log and 10 will cancel each other out) x= 100- 9 x=91 IF u have any question please ask, or need more explanation :D